OpenStudy (anonymous):

What logarithmic function represents the data in the table? x f(x) 216 3 1296 4 7776 5

OpenStudy (anonymous):

these are thee answers f(x) = logx 5 f(x) = 6 log x f(x) = log 6x f(x) = log6 x

OpenStudy (anonymous):

Can you rewrite your answer possibilities with the equation tool so we can completely understand. There are some ambiguities.

OpenStudy (anonymous):

no there is no power signs or anything its correct for sure.

OpenStudy (anonymous):

Are the log's all base ten, besides the last one?

OpenStudy (anonymous):

wait i think i found out what you ment here the number closer to the log sign is a base and the farther one is just regular

OpenStudy (anonymous):

Do me a solid and just rewrite them using the tool.

OpenStudy (anonymous):

\[\log _{x}5\] \[6 \log x\] \[\log 6^{x}\] \[\log _{6}x\]

OpenStudy (anonymous):

Thank you. All you need to do is guess and check. Plug in the value of x for each possibility and see if it gives you the f(x) value you're looking for.

OpenStudy (anonymous):

This is how you solve logarithms of any base. http://www.purplemath.com/modules/logrules5.htm This, and a calculator, will do the trick.